3. A group can charter an aircraft at a fixed total cost. If 36 ppl charter the aircraft rather than 40, then the cost per person is greater by $12. What is the cost per person if 40 people charter the aircraft?



Answer by stanbon(75887)

A group can charter an aircraft at a fixed total cost. If 36 ppl charter the aircraft rather than 40, then the cost per person is greater by $12. What is the cost per person if 40 people charter the aircraft?
---------------------------------
Let the fixed cost be "x" dollars
-------------------
1st average cost: x/40 dollars
--------------
2nd average cost; x/36 dollars
----------------
EQUATION:
36 person cost - 40 person cost = 12 dollars
x/36 - x/40 = 12
(40x-36x) = 12*36*40
4x = 12*36*40
x = 3*36*40
x = $4320
x/40 = $108 (average cost for 40 persons)
====================
Cheers,
Stan H.